OpenStudy (anonymous):

Find the number of real solutions for the equation x^2-2x+3=0

OpenStudy (john_es):

If you can use the quadratic formula, \[ax^2+bx+c=0\\ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{2\pm\sqrt{4-12}}{2}\] It has no real solutions.